# Eigen-Value: Efficient Domain-Robust Data Valuation via Eigenvalue-Based Approach

### v2

@@ Abstract
Data valuation has become central in the era of data-centric AI. It drives efficient training pipelines and enables objective pricing in data markets by assigning a numeric value to each data point. Most existing data valuation methods estimate the effect of removing individual data points by evaluating changes in model validation performance under in-distribution (ID) settings, as opposed to out-of-distribution (OOD) scenarios where data follow different patterns. Since ID and OOD data behave differently, data valuation methods based on ID loss often fail to generalize to OOD settings, particularly when the validation set contains no OOD data. Furthermore, although OOD-aware methods exist, they involve heavy computational costs, which hinder practical deployment. To address these challenges, we introduce Eigen-Value (EV), a plug-and-play data valuation framework for OOD robustness that uses only an ID data subset, including during validation. EV provides a new spectral approximation of domain discrepancy, which is the gap of loss between ID and OOD using ratios of eigenvalues of ID data’s covariance matrix. EV then estimates the marginal contribution of each data point to this discrepancy via perturbation theory, alleviating the computational burden. Subsequently, EV plugs into ID loss-based methods by adding an EV term without any additional training loop. We demonstrate that EV achieves improved OOD robustness and stable value rankings across real-world datasets, while remaining computationally lightweight. These results indicate that EV is practical for large-scale settings with domain shift, offering an efficient path to OOD-robust data valuation.
@@ 1 Introduction
Machine learning has achieved strong performance in image recognition, autonomous driving, and conversational systems. Yet domain shifts between in-distribution (ID) training data and out-of-distribution (OOD) deployment data can sharply reduce accuracy. Robustness to such shifts is essential, especially in safety-critical applications. Most work addresses this with model-centric strategies such as distributionally robust optimization Hu et al. [2018], Staib and Jegelka [2019], Rahimian and Mehrotra [2022] and domain-invariant representation learning, which often require specialized architectures and complex training pipelines. However, even with state-of-the-art architectures and training algorithms, performance remains bounded by data quality and composition. A complementary data-centric approach called data valuation evaluates the training set itself and identifies informative examples Sim et al. [2022], Tian et al. [2022], Agarwal et al. [2019]. Curating data, rather than continually updating models, offers multiple benefits. It reduces computation, accelerates learning, and improves model performance Xu et al. [2025]. This view also aligns with pricing in data marketplaces.
Data valuation measures the change in model performance induced by including or excluding each training example. Most existing methods estimate this effect under the training distribution Shapley et al. [1953], Roth [1988], Ghorbani and Zou [2019], Alvarez-Melis and Fusi [2020]. However, validation and OOD distributions often diverge, so values inferred from validation-based performance changes do not hold under OOD. This gap in distributional alignment makes OOD-robust data valuation necessary. However, existing shift-aware methods are computationally prohibitive Lin et al. [2024], limiting their practicality in data marketplaces that require reliable metrics computed without OOD data Sim et al. [2022], Tian et al. [2022].
We address this gap with Eigen-Value (EV), a data valuation framework that targets robustness under domain shift. First, EV establishes a connection between domain discrepancy and the Hessian of the loss function under distribution shift. This relies on the observation that, in logistic regression, the Hessian approximates the data covariance Le Cun et al. [1991], Lin et al. [2007], Hazan et al. [2014]. Building on this observation, EV introduces a novel formulation that relates domain discrepancy to the eigenvalues of the covariance structure, as shown in Figure 1. Second, to address the computational burden of repeated eigendecomposition, EV employs perturbation theory Kato [2013], which approximates the effect of removing a single data point without requiring a new decomposition. Lastly, EV augments existing ID loss-based data valuation methods with a marginal value for the domain discrepancy term that reflects the induced eigenvalue shifts. Importantly, EV operates solely on ID data, requiring no OOD samples for training or validation. It estimates the potential impact of domain shifts by quantifying how each data point perturbs the largest and smallest eigenvalues in the ID setting. In experiments across diverse datasets, augmenting baselines with EV improves OOD performance over the baselines alone, while maintaining efficiency and stability.
@@
    We relate domain discrepancy to covariance eigenvalues, which enables data valuation without OOD samples.
    We introduce EV, a scalable and easy-to-combine term that upgrades ID-based methods via perturbation theory.
    We present evidence on real-world datasets that EV improves OOD robustness, stability, and efficiency, indicating readiness for practical use.
@@ 2.1 Data Valuation
Data valuation measures how each training example changes a model’s performance Sim et al. [2022], Sidi et al. [2012]. Its importance grows with large-scale datasets and retrieval augmented generation Lewis et al. [2020], where careful curation improves efficiency and interpretability Koh and Liang [2017]. Data Shapley Ghorbani and Zou [2019] estimates a point’s value by retraining across many subsets, which is expensive. KNN Shapley Jia et al. [2019] and Data-OOB Kwon and Zou [2023] reduce cost using local neighbors and out-of-bag scores, but they face limits in scalability or accuracy. LAVA Just et al. [2023] removes the retraining bottleneck by measuring Wasserstein distances and gives fast estimates. However, its effectiveness is limited to data sampled from the same distribution as the training set, leading to degraded utility under domain shift. Deviation Lin et al. [2024] formulates a worst-case distributional shift objective via the neural tangent kernel (NTK), enabling applicability beyond the training distribution. However, it requires $n$ times of inverting an $n\times n$ kernel matrix (with $n$ training samples), incurring huge computation, which hampers practical deployment at scale. Even on reduced subsets, the method can be unstable due to its worst-case formulation Zhai et al. [2021], where small changes in the dataset lead to significant fluctuations in value rankings, limiting its suitability for data-centric AI at scale. Our work departs from these existing methods and proposes an eigenvalue-based scheme that remains accurate under domain shift and is accelerated by perturbation theory.
@@ 2.2 OOD Robustness
OOD robustness seeks models that remain reliable when the test distribution differs from that of the training and validation data  Yadav et al. [2023], Oh et al. [2024], Wortsman et al. [2022]. Distribution shifts, from small corruptions to large domain gaps, can reduce accuracy even for state-of-the-art networks, so training and evaluation must anticipate such changes. Recent advances include ensembling and spectral criteria. WiSE-FT Wortsman et al. [2022] ensembles pre- and post-fine-tuned weights to improve both ID and OOD performance. RankMe Garrido et al. [2023] and CaRoT Oh et al. [2024] relate generalization to spectra of weight or feature matrices and guide optimization toward high rank or large minimum singular values. These results suggest that spectral or weight space signals are useful proxies for OOD risk. However, most of this literature on OOD robustness has been model-centric, and there has been a lack of an efficient data valuation method for assessing OOD loss. To enable efficient data pipeline management and provide objective pricing in data marketplaces, we propose an OOD-robust data valuation method. This data-centric approach allows us to evaluate OOD robustness from the perspective of data itself.
@@ 2.3 Eigenvalue Methods
Eigenvalue analysis has guided machine learning since classical PCA Hotelling [1933]. Its view of covariance spectra also influenced kernel PCA, spectral clustering, and diffusion maps Maćkiewicz and Ratajczak [1993], Schölkopf et al. [1997], Von Luxburg [2007], Lafon [2004]. Recent works use spectral information to improve robustness. CaRoT Oh et al. [2024] steers parameter updates using Hessian eigenvalues, and Eigen-SAM Luo et al. [2024] discourages sharp directions to find flatter minima. We extend this spectral line to data valuation. By linking covariance eigenvalues to OOD generalization error and using perturbation theory to approximate each sample’s spectral influence, our method, EV provides a scalable valuation approach aware of domain shift.
@@ 3 Setting and Preliminaries
Figure 2: PCA visualization of normalized embeddings sampled (1K each) from different domain sources. The two distributions, corresponding to different domains, partially overlap due to normalization, illustrating that the matching marginal assumption remains applicable in real-world scenarios.
@@
Let $S=\{(x_{i},y_{i})\}_{i=1}^{N}$ be a dataset of NN normalized (embedding, label) pairs with $x_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ and $y_{i}\in\mathcal{Y}$. For any subset $\mathcal{D}\subseteq S$ of size $n$, a utility function $U:2^{n}\!\rightarrow\!\mathbb{R}$ maps the subset to a scalar score. We define the marginal data value $V:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ of a point $(x_{k},y_{k})$ as the drop in utility when that point is removed, i.e., $V(x_{k},y_{k})\;=\;U(\mathcal{D})\;-\;U\!\bigl(\mathcal{D}\setminus\{(x_{k},y_{k})\}\bigr)$. Thus $V(x_{k},y_{k})>0$ indicates that deleting the point hurts performance, implying high importance. Existing methods share this core definition and differ only in how they approximate $U$ and compute $V$ efficiently.
@@
Let $P_{\text{ID}}(x)$ denote the in-distribution (ID) used for training, and $P_{\text{OOD}}(x)$ an out-of-distribution (OOD) that differs only in the marginal $P(x)$ but shares the same conditional $P(y\mid x)$ Shimodaira [2000], Moreno-Torres et al. [2012]. In other words, the input distribution shifts between ID and OOD, yet the relationship between input and label remains identical Sugiyama et al. [2007]. To further analyze this OOD scenario, we consider the matching marginal condition. This condition implies that the covariance matrices of ID and OOD share the same diagonal entries, meaning that each feature has identical variance across domains. In contrast, the off-diagonal elements may differ Huang et al. [2006], reflecting possible shifts in cross-feature correlations. Although normalization is not inherent to the definition, in our setting, it makes the condition easier to satisfy by aligning the diagonal entries. As illustrated in Figure 2, PCA visualizations of 1K embeddings sampled from different domains show partial overlap after normalization, supporting that the matching marginal assumption is both empirically plausible and practically valid.
@@
Perturbation theory offers a mathematical framework for approximating how a matrix’s eigenvalues and eigenvectors shift under a perturbation. For instance, if a matrix AA is modified by a small term $\epsilon$ such that $B=A+\epsilon$, one can express the eigenvalues and eigenvectors of $B$ as expansions in $\epsilon$ Moro and Dopico [2003], Greenbaum et al. [2020]. In the context of covariance matrices, this technique elucidates how adjustments to the underlying data affect eigenvalue structures Sugiyama et al. [2020], Mohammed et al. [2017]. Consequently, removing a single data point $k$ from a dataset can be viewed as applying a perturbation to the covariance matrix, thereby enabling an analysis of how individual samples influence eigenvalue shifts, as shown in Section 4.2. This perspective aids in evaluating dataset stability and identifying observations that exert outsized influence on the learned model.
@@ 4 Efficient Domain Robust Data Valuation
We propose an eigenvalue-based framework for OOD-robust data valuation, ensuring reliability under distribution shifts. First, we characterize domain discrepancy using eigenvalues of covariance matrix derived from normalized training dataset with zero mean $\mathcal{D}_{\text{ID}}=\{z_{i}=(x_{i},y_{i})\}_{i=1}^{n}$, which is i.i.d. sampled from the ID distribution to quantify shifts between ID and OOD data (Section 4.1). After that, we develop an efficient perturbation-based method to compute marginal valuations directly from eigenvalue terms, reducing computational overhead (Section 4.2). Finally, we integrate these components into a unified valuation framework, enabling scalable and robust data valuation without requiring explicit OOD samples (Section 4.3).
@@ 4.1 Utility Function Based on OOD Loss
In the Shapley value framework, the utility function $U$, which forms the basis for data valuation, is typically defined as the model’s performance on a validation set, with $U(\mathcal{D})$ denoting validation performance on dataset $\mathcal{D}$. Similarly, we adopt $\mathcal{L}_{\text{OOD}}(\theta)$ as our utility function, which is the loss function of a model with parameter $\theta$ trained on the ID set $\mathcal{D}_{\text{ID}}$ and evaluated on OOD set $\mathcal{D}_{\text{OOD}}$. The OOD loss $\mathcal{L}_{\text{OOD}}(\theta)$ can be upper-bounded using the ID loss $\mathcal{L}_{\text{ID}}(\theta)$ and a measure of domain discrepancy $\Gamma(\mathcal{D}_{\text{OOD}},\mathcal{D}_{\text{ID}}):=\sup_{\theta}\left|\mathcal{L}_{\text{OOD}}(\theta)-\mathcal{L}_{\text{ID}}(\theta)\right|$. For notational simplicity, we denote $\mathcal{L}_{\text{OOD}}(\theta)$ and $\mathcal{L}_{\text{ID}}(\theta)$ as $\mathcal{L}_{\text{OOD}}$ and $\mathcal{L}_{\text{ID}}$.
@@
To quantify domain discrepancy, we first define it in terms of the distributional shift between ID and OOD data. A well-established approach is to approximate domain shift using a measure derived from the model’s sensitivity to input variations. Notably, when the loss function is formulated as Normalized Cross Entropy (NCE) Ma et al. [2020], domain discrepancy can be related to the spectral properties of the Hessian matrices $H_{\text{ID}}=\nabla^{2}_{\theta}\mathcal{L}_{\text{ID}}$ and $H_{\text{OOD}}=\nabla^{2}_{\theta}\mathcal{L}_{\text{OOD}}$, corresponding to the ID and OOD distributions, respectively.
Proposition 1.
@@
Theorem 1.
We assume $\Sigma_{\textnormal{OOD}}=\Sigma_{\textnormal{ID}}+E$, where $E\in\mathbb{R}^{d\times d}$ has zero diagonal entries and non-zero off-diagonal elements representing domain discrepancies. Based on this assumption, we derive the following bound on $\mathcal{L}{\textnormal{OOD}}$ in terms of $\lambda{\max}$, $\lambda_{\min}$, and the dimensionality $d$ of $\Sigma_{\textnormal{ID}}$.
@@ 4.2 Marginal Calculation of Eigenvalue term
We denote $\delta_{\max}^{(k)}:=u_{\max}^{\top}\Delta_{k}u_{\max}$, $\delta_{\min}^{(k)}:=u_{\min}^{\top}\Delta_{k}u_{\min}$, capturing the sensitivity of the eigenvalues to the removal of data point $x_{k}$ in Eq. 5. To evaluate how well this approximation estimates the eigenvalue shift, we conducted an experiment using each 1K normalized embedding data point. Specifically, we examined the actual difference in eigenvalues with and without $x_{k}$ and assessed the linearity of our proposed approximation. Figure 3 demonstrates a consistent linear relationship between the actual eigenvalue difference and the estimated value. We observed that the difference in eigenvalues can be approximated using our proposed method.
Leveraging this insight, we approximate the difference in the eigenvalue term from Eq. 3 by directly substituting the approximation of the difference in eigenvalues. More concretely, when considering the covariance matrix without data instance $k$, we replace its eigenvalue with $\delta_{\max}^{(k)},\delta_{\min}^{(k)}$. Then, by applying a Taylor expansion, we can use the resulting approximated terms, together with the eigenvalues of the full data covariance matrix, to approximate the marginal value of data point $x_{k}$ to the domain discrepancy.
Figure 3: Relation between approximation values $u^{\top}_{\text{max}}\Delta_{k}u_{\text{max}}$ in Eq. 5 and real values $\lambda_{\text{max}}(\Sigma_{-k})-\lambda_{\text{max}}(\Sigma_{\textnormal{ID}})$ for CIFAR-10, ImageNet, Amazon Reviews - Books and DomainNet - Real embedding datasets. This demonstrates that eigenvalue differences can be accurately approximated using our proposed method, highlighting its effectiveness in capturing spectral variations.
Theorem 2.
@@
By leveraging this approximation, we can estimate the marginal value of each data point using perturbation theory without having to repeatedly perform costly eigendecompositions. This makes OOD-robust data valuation computationally efficient and feasible for practical deployment.

### v3

@@ Abstract
Data valuation is central to data-centric AI, enabling efficient training and objective data pricing by assigning a value to each data point. Most existing methods estimate the effect of removing data points based on in-distribution (ID) validation performance, often failing to generalize to out-of-distribution (OOD) settings where data follow different patterns. Although OOD-aware methods exist, they incur substantial computational costs. To address this, we propose Eigen-Value (EV), a plug-and-play data valuation framework for OOD robustness using only ID data. EV approximates domain discrepancy, defined as the gap between ID and OOD loss, via eigenvalue ratios of the ID covariance matrix, and estimates each data point’s marginal contribution using perturbation theory. It integrates seamlessly with ID loss–based methods without additional training. Experiments show that EV improves OOD robustness and ranking stability across real-world datasets while remaining computationally efficient, making it practical for large-scale data valuation under domain shift.
@@ 1 Introduction
Machine learning has achieved remarkable success in image recognition, autonomous driving, and conversational systems. However, domain shifts between in-distribution (ID) training data and out-of-distribution (OOD) deployment data can drastically degrade performance, posing risks in safety-critical settings. Most studies address this challenge through model-centric strategies such as distributionally robust optimization [17, 48, 40] or domain-invariant representation learning, often relying on specialized architectures and complex pipelines. A complementary data-centric paradigm, data valuation, identifies informative examples [47, 51, 1]. Curating data rather than continually updating models reduces computation, accelerates learning, and enhances performance [57], while aligning with data marketplaces’ pricing mechanisms.
Data valuation measures the change in model performance induced by including or excluding each training example. Most existing methods estimate this effect under the training distribution [45, 42, 8]. However, validation and OOD distributions often diverge, so values inferred from validation-based performance changes do not hold under OOD. This gap in distributional alignment makes OOD-robust data valuation necessary. However, existing shift-aware methods are computationally prohibitive [30], limiting their practicality in data marketplaces that require reliable metrics computed without OOD data [47, 51].
We address this gap with Eigen-Value (EV), a data valuation framework for robustness under domain shift. EV connects domain discrepancy to the Hessian of the loss function, where in logistic regression the Hessian approximates data covariance [27, 29, 10]. Based on this, EV formulates domain discrepancy via eigenvalues of the covariance structure (Figure 1). To reduce the cost of repeated eigendecomposition, it applies perturbation theory [22] to approximate single-removal efficiently. EV extends ID loss–based valuation methods by incorporating a marginal value for eigenvalue shifts induced by domain discrepancy. Using ID data, EV estimates OOD robustness by quantifying how each sample perturbs the largest and smallest eigenvalues. Experiments show that augmenting baselines with EV improves OOD performance while maintaining efficiency.
@@
  • We relate domain discrepancy to covariance eigenvalues, enabling data valuation without OOD samples.
  • We introduce EV, a scalable and easily integrable term that upgrades ID-based methods via perturbation theory.
  • We show on real-world datasets that EV improves OOD robustness, stability, and efficiency.
@@ 2.1 Data Valuation
Data valuation measures how each training example changes a model’s performance [47]. Its importance grows with large-scale datasets and retrieval augmented generation [28], where careful curation improves interpretability [23]. Data Shapley [8] estimates a point’s value by retraining across many subsets, which is expensive. KNN Shapley [19] and Data-OOB [25] reduce cost using local neighbors and out-of-bag scores, but face limits in scalability or accuracy. LAVA [21] removes the retraining bottleneck by measuring Wasserstein distances and gives fast estimates. However, its effectiveness is limited to data sampled from the same source distribution as the model’s training set, leading to degraded utility under domain shift. Deviation [30] formulates a worst-case distributional shift objective via the neural tangent kernel (NTK), enabling applicability beyond the training distribution. However, it requires $n$ inversions of $n\times n$ kernel matrix (with $n$ samples), incurring huge computation that hampers practical deployment at scale. Even on reduced subsets, the method may exhibit instability, as its worst-case formulation [59] can amplify sensitivity to small changes, leading to fluctuations in value rankings and limiting suitability for data-centric AI at scale. Our work departs from these existing methods and proposes an eigenvalue-based scheme that remains accurate under domain shift and is accelerated by perturbation theory.
@@ 2.2 OOD Robustness
OOD robustness seeks models that remain reliable when the test distribution deviates from that of the training and validation data  [38, 55]. Distribution shifts, from small corruptions to large domain gaps, can reduce accuracy even for state-of-the-art networks, so training and evaluation must anticipate such changes. Recent advances include ensembling and spectral criteria. WiSE-FT [55] ensembles pre- and fine-tuned weights to improve both ID and OOD performance. RankMe [7] and CaRoT [38] relate generalization to spectra of weight or feature matrices and guide optimization toward high rank or large minimum singular values. These results suggest that spectral or weight space signals are useful proxies for OOD risk. However, most of this literature on OOD robustness has been model-centric, and there has been a lack of an efficient data valuation method for assessing OOD loss. To enable efficient data pipeline management and provide objective pricing in data marketplaces, we propose an OOD-robust data valuation method. This data-centric approach allows us to evaluate OOD robustness from the perspective of data itself.
@@ 2.3 Eigenvalue Methods
Eigenvalue analysis has guided machine learning since classical PCA [15]. Its view of covariance spectra also influenced kernel PCA, spectral clustering, and diffusion maps [34, 44, 52, 26]. Recent works use spectral information to improve robustness. CaRoT [38] encourages a larger smallest singular value of the ID input covariance matrix, and Eigen-SAM [32] discourages sharp directions to find flatter minima. We extend this spectral line to data valuation. By linking covariance eigenvalues to OOD generalization error and using perturbation theory to approximate each sample’s spectral influence, our method EV provides a scalable valuation approach aware of domain shift.
@@ 3 Setting and Preliminaries
Let $S=\{(x_{i},y_{i})\}_{i=1}^{N}$ be a dataset of $N$ normalized (embedding, label) pairs with $x_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ and $y_{i}\in\mathcal{Y}$. For any subset $\mathcal{D}\subseteq S$ of size $n$, a utility function $U:2^{n}\!\rightarrow\!\mathbb{R}$ maps the subset to a scalar score. We define the marginal data value $V:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ of a point $(x_{k},y_{k})$ as the drop in utility when that point is removed, i.e., $V(x_{k},y_{k})\;=\;U(\mathcal{D})\;-\;U\!\bigl(\mathcal{D}\setminus\{(x_{k},y_{k})\}\bigr)$. Thus $V(x_{k},y_{k})>0$ indicates that deleting the point hurts performance, implying high importance. Existing methods share this core definition and differ only in how they approximate $U$ and compute $V$ efficiently.
@@
Let $P_{\text{ID}}(x)$ denote the in-distribution (ID) used for training, and $P_{\text{OOD}}(x)$ an out-of-distribution (OOD) that differs only in the marginal $P(x)$ but shares the same conditional $P(y\mid x)$ [46, 36]. In other words, the input distribution shifts between ID and OOD, yet the relationship between input and label remains identical [49]. To further analyze this OOD scenario, we consider the matching marginal condition. This condition implies that the covariance matrices of ID and OOD share the same diagonal entries, meaning that each feature has identical variance across domains. In contrast, the off-diagonal elements may differ [18], reflecting possible shifts in cross-feature correlations. Although normalization is not inherent to the definition, we apply feature-wise normalization to both ID and OOD data in our setting. This step aligns the diagonal entries of their covariance matrices and makes the condition easier to satisfy.
@@
Perturbation theory offers a mathematical framework for approximating how a matrix’s eigenvalues and eigenvectors shift under a perturbation. For instance, if a matrix $A$ is modified by a small term $\epsilon$ such that $B=A+\epsilon$, one can express the eigenvalues and eigenvectors of $B$ as expansions in $\epsilon$ [37, 9]. In the context of covariance matrices, this technique elucidates how adjustments to the underlying data affect eigenvalue structures [50, 35]. Consequently, removing a single data point $k$ from a dataset can be viewed as applying a perturbation to the covariance matrix, thereby enabling an analysis of how individual samples influence eigenvalue shifts, as shown in Section 4.2. This perspective aids in evaluating dataset stability and identifying observations that exert outsized influence on the learned model.
@@ 4 Efficient Domain Robust Data Valuation
We propose an eigenvalue-based framework for OOD-robust data valuation, focusing on the covariate-shift setting to ensure reliability under distribution shifts. First, we characterize domain discrepancy using eigenvalues of covariance matrix derived from normalized training dataset with zero mean $\mathcal{D}_{\text{ID}}=\{z_{i}=(x_{i},y_{i})\}_{i=1}^{n}$, which is i.i.d. sampled from the ID distribution to quantify shifts between ID and OOD data (Section 4.1). After that, we develop an efficient perturbation-based method to compute marginal valuations directly from eigenvalue terms, reducing computational overhead (Section 4.2). Finally, we integrate these components into a unified valuation framework, enabling scalable and robust data valuation without requiring explicit OOD samples (Section 4.3).
@@ 4.1 Utility Function Based on OOD Loss
In the Shapley value framework, the utility function $U$, which forms the basis for data valuation, is typically defined as the model’s results on a validation set, with $U(\mathcal{D})$ denoting validation performance on dataset $\mathcal{D}$. Similarly, we adopt $\mathcal{L}_{\text{OOD}}(\theta)$ as our utility function, which is the loss function of a model with parameter $\theta$ trained on the ID set $\mathcal{D}_{\text{ID}}$ and evaluated on OOD set $\mathcal{D}_{\text{OOD}}$. The OOD loss $\mathcal{L}_{\text{OOD}}(\theta)$ can be upper-bounded using the ID loss $\mathcal{L}_{\text{ID}}(\theta)$ and a measure of domain discrepancy $\Gamma(\mathcal{D}_{\text{OOD}},\mathcal{D}_{\text{ID}}):=\sup_{\theta}\left|\mathcal{L}_{\text{OOD}}(\theta)-\mathcal{L}_{\text{ID}}(\theta)\right|$. For notational simplicity, we denote $\mathcal{L}_{\text{OOD}}(\theta)$ and $\mathcal{L}_{\text{ID}}(\theta)$ as $\mathcal{L}_{\text{OOD}}$ and $\mathcal{L}_{\text{ID}}$.
@@
To quantify domain discrepancy, we first define it in terms of the distributional shift between ID and OOD data. A well-established approach is to approximate domain shift using a measure derived from the model’s sensitivity to input variations. Notably, when the loss function is formulated as Normalized Cross Entropy (NCE) [33], domain discrepancy can be related to the spectral properties of the Hessian matrices $H_{\text{ID}}=\nabla^{2}_{\theta}\mathcal{L}_{\text{ID}}$ and $H_{\text{OOD}}=\nabla^{2}_{\theta}\mathcal{L}_{\text{OOD}}$, corresponding to the ID and OOD distributions, respectively.
Proposition 1
@@
Theorem 1
We assume $\Sigma_{\textnormal{OOD}}=\Sigma_{\textnormal{ID}}+E$, where $E\in\mathbb{R}^{d\times d}$ has zero diagonal entries and non-zero off-diagonal elements representing domain discrepancies. Based on this assumption, we derive the following bound on $\mathcal{L}_{\textnormal{OOD}}$ in terms of $\lambda_{\max}$, $\lambda_{\min}$, and the dimensionality $d$ of $\Sigma_{\textnormal{ID}}$.
@@ 4.2 Marginal Calculation of Eigenvalue term
We denote $\delta_{\max}^{(k)}:=u_{\max}^{\top}\Delta_{k}u_{\max}$, $\delta_{\min}^{(k)}:=u_{\min}^{\top}\Delta_{k}u_{\min}$, capturing the sensitivity of the eigenvalues to the removal of data point $x_{k}$ in Eq. 5. To evaluate how well this approximation estimates the eigenvalue shift, we conducted an experiment using each 1k normalized embedding data point. Specifically, we examined the actual difference in eigenvalues with and without $x_{k}$ and assessed the linearity of our proposed approximation. Figure 2 shows a clear linear trend between the true and estimated eigenvalue differences, supported by a low RMSE and a high Pearson correlation. We observed that the difference in eigenvalues can be approximated using our proposed method.
Leveraging this insight, we approximate the difference in the eigenvalue term from Eq. 3 by directly substituting the approximation of the difference in eigenvalues. More concretely, when considering the covariance matrix without data instance $k$, we replace its eigenvalue with $\delta_{\max}^{(k)},\delta_{\min}^{(k)}$. Then, by applying a Taylor expansion, we can use the resulting approximated terms, together with the eigenvalues of the covariance matrix, to approximate the marginal value of data point $x_{k}$ to the domain discrepancy.
Figure 2: Relation between approximation values $u^{\top}_{\text{max}}\Delta_{k}u_{\text{max}}$ in Eq. 5 and real values $\lambda_{\text{max}}(\Sigma_{-k})-\lambda_{\text{max}}(\Sigma_{\textnormal{ID}})$ for CIFAR-10, ImageNet, Amazon Reviews - Books and DomainNet - Real embedding datasets. This demonstrates that eigenvalue differences can be accurately approximated using our method. The low RMSE and high Pearson correlation demonstrate the reliability of the approximation.
Theorem 2
@@
By leveraging this approximation, we can quantify the marginal value of each data point using perturbation theory without repeatedly performing costly eigendecompositions. This makes OOD-robust data valuation computationally efficient and feasible for practical deployment, achieving a complexity of $O(nd^{2}+d^{3})$, which is highly efficient in regimes where $n\gg d$.